# Federated attention contrastive learning models for prostate cancer diagnosis and Gleason grading

### v4

@@ 3 Methods
In this section, we provide an overview of the proposed FACL framework, as illustrated in Fig. 1. FL retains data locally on individual centers, allowing for patch-level segmentation and feature extraction of tissue regions directly on pathological images at the client level. In each client site, WSIs are cropped into patch images with a size of $224\times 224$ pixels. Then, the Swin Transformer, augmented with convolutional operations, is employed to extract discriminative features. These features are then locally preserved as input for subsequent model training within the federated learning framework. Each client trains the model by transmitting model parameters to the central server. Subsequently, the server performs a weighted averaging upon collecting all client model parameters and then distributes the updated parameters to all centers for the next round of training. In addition to these procedures, we enhance dataset protection by incorporating differential privacy. This involves introducing random noise during the aggregation of model parameters at the server. Finally, to address the problem of unbalanced data distribution among centers, we also use the attention comparison to improve the consistency between the server and client models, which results in an improved final training model.
Initially, each client employs the CTransPath feature extractor wang2022transformer to acquire and store features from their local datasets. Subsequently, within the FACL framework, we integrate AttMIL ilse2018attention , a weakly supervised model with an attention mechanism, for diagnosing and grading gigapixel WSIs with prostate cancer. This approach not only ensures patient privacy but also boosts the model’s performance trained on individual client data. To tackle the Non-IID challenge, we introduce attention contrastive learning. Despite imbalanced data distribution among local centers, FACL can rectify these deviations, enabling the model to achieve optimal results.
@@ 3.1 Data preprocessing
High-resolution histopathology images often demand time-consuming analysis and are susceptible to interference from complex backgrounds wang2023sac . An integral step in data processing involves cropping WSIs into smaller sizes suitable for model input. Removing the background region is crucial since it lacks informative content. We cropped patches ($224\times 224$ pixels) without overlap in regions with significant tissue presence, utilizing the Otsu thresholding method 4310076 . Subsequently, the features of each patch were extracted using the CTransPath feature extractor wang2022transformer , and these features were aggregated for consolidated storage across the entire image. It is noteworthy that, unlike the ImageNet pre-trained model, the CTransPath extractor is pre-trained on multiple histopathology datasets.
@@ 3.2 Weakly-supervised learning on WSIs
At each local center, we employ a weakly supervised model, AttMIL, detailed in ilse2018attention . AttMIL integrates an attention mechanism, this model forms the basis for FL across local centers. Fig. 2 depicts the attention mechanism pipeline, comprising three modules: $f_{pro}$ for projection, $f_{att}$ for attention, and $f_{pre}$ for prediction. We utilize the gating mechanism in conjunction with the $\tanh(\cdot)$ non-linearity. This mechanism enables the model to dynamically modulate the importance of different parts of the input sequence when making predictions. In this context, we utilize the feature vector extracted from CTransPath as the model’s input. The attention mechanism facilitates the computation of attention scores for each instance, aiding in the identification of similarities or dissimilarities between instances. This approach, particularly relevant in computational pathology, provides interpretability and reasoning for regions of interest. This is particularly beneficial for medical professionals in the diagnosis of WSIs.
Fig. 2: The specific implementation details of the attention mechanism in FACL.
We provide detailed layer-by-layer explanations beginning with the projection module. This module, comprising contiguous, trainable, fully-connected layers, transforms fixed feature embeddings from a pre-trained encoder into a more compact feature space tailored to histopathology images related to the chosen disease model. The initial fully-connected layer $\bm{W_{1}}\in\mathbb{R}^{512\times 768}$ compresses the 768-dimensional patch-level features $\bm{z}_{k}$ into a 512-dimensional vector $\bm{h}_{k}=\bm{W}_{1}\bm{z}{k}^{\top}$. The resulting set of $N$ patch-level $\bm{h}_{k}$ in a whole slide image is represented by $\bm{H}\in\mathbb{R}^{N\times 512}$.
If we consider the first two attention matrices, $\bm{U}a$ and $\bm{V}a$, both of dimensions $\mathbb{R}^{256\times 512}$, as the shared attention backbone for all classes, the attention network can be divided into n parallel attention branches, denoted as $\bm{W}_{a,1},...,\bm{W}_{a,n}$, each with dimensions $\mathbb{R}^{1\times 256}$. Simultaneously, create $n$ parallel independent classifiers, $\bm{W}_{c,1},...,\bm{W}_{c,n}$, to assess each class-specific slide-level representation. The attention score for the kth patch of the mth class denoted as $a_{k,m}$ and computed by equation (1), contributes to the aggregation of the slide-level feature representation. The resulting slide-level feature representation for the mth class is denoted as $h_{slide,m}\in\mathbb{R}^{1\times 512}$ and is computed by equation (2).
@@
The slide-level score $s_{slide,m}$ is derived from the classification layer $\bm{W}_{c,m}\in\mathbb{R}^{1\times 512}$ through the formula $s_{slide,m}=\bm{W}_{c,m}h_{slide,m}^{\top}$. During inference, the predicted probability distribution for each class is obtained by applying the softmax function to the slide-level prediction scores.
@@ 3.3 Federated learning on WSIs
In this section, we elaborate on the details of the proposed federated attention contrastive learning method, as well as the differential privacy techniques used to provide additional protection to the model.
3.3.1 Federated attention contrastive learning
Fig. 3: Illustration of the difference between conventional contrastive learning and attention contrastive learning. $P(x)$ in (a) represents the representation of $x$. $A(x)$ in (b) represents the attention obtained by $x$ through model $M$ ($M_{l}$ and $M_{s}$ correspond to the client model and server model, respectively).
Inspired by contrastive learning, we considered applying it at the attention level. Fig. 3 illustrates the fundamental difference between conventional contrastive learning and attention contrastive learning methods. In conventional contrastive learning, the model is trained to maximize the proximity of similar samples while maintaining a separation from dissimilar ones. A key feature of this learning method is that it does not require explicit label usage in the training process. In our attention contrastive learning, we integrate the characteristics of client and server models in the FL learning, aiming to maximize the consistency between the attention scores of the local client model and that of the server model. This ensures that a center model indirectly acquires the attentional insights of other centers’ models through the intermediary server. Based on attention contrastive learning, we propose FACL by combining the FedAvg mcmahan2017communication and weakly supervised learning. It introduces lightweight and efficient modifications in client-side training. Each client trains a relatively personalized model during local training due to data heterogeneity. The server’s global model, in turn, acquires a significantly improved representation compared to the local model. The aim of FACL is to bolster the learning and generalization capabilities of the client-side model by augmenting the consistency between the attention learned by the client-side model and that acquired by the server-side model.
The loss of the local model consists of two components. The first part is the standard loss, such as cross-entropy loss in supervised learning, denoted as $\mathcal{L}_{CE}$ in this paper. The second part is the contrastive loss, denoted as $\mathcal{L}_{FACL}$, which measures the disparity between the attention scores of client-side models and the server-side model. Given the predicted probability value of the classification model as ”logits,” the label as ”y,” and the attention scores of the local model and the server model as $A_{local}$ and $A_{s}$, respectively, the total loss can be expressed as follows:
@@ 10                   $s_{i,j}=f_{pre}(\bm{h}_{bag_{i,j}})$ ;
                   // Calculate attention similarity
@@ 18             $\bm{W}_{i}\leftarrow\bm{W}_{s}$ ;
Here, $\bm{\mu}$ represents the hyperparameter controlling the weight of the attention contrastive loss. $\mathcal{L}_{FACL}$ denotes the Kullback-Leibler divergence joyce2011kullback , quantifying the similarity between two probability distributions.
@@ 3.3.2 FACL with differential privacy
By embracing the principles of data differential privacy, we seamlessly integrated this approach into the FL server, specifically during the phase where the server receives parameter updates from individual client models. The crux of differential privacy involves introducing noise to the model weights before transmitting them to the centralized server. To execute this, we utilized a noise generator denoted as $N_{g}(\cdot)$, generating Gaussian random noise $\bm{Z}\sim(0,\bm{z}^{2}\eta^{2})$, with $\bm{z}$ denoting the noise level and $\eta$ representing the standard deviation of the neural network weight.
The noise is carefully calibrated to prevent significant degradation of learned weights or overall model performance. Striking the right balance involves introducing sufficient noise to safeguard privacy while enabling the model to extract valuable information adnan2022federated ; ziller2021medical . In FL, noisy weight updates from multiple centers are consolidated on the server. The noise introduced by each center tends to offset when combined, resulting in a more precise global model. This occurs because the noise is typically drawn from a distribution with a mean of zero, leading to an averaging effect when aggregated from different centers. Although the introduction of noise in the weights may marginally influence learned weights, the FL is designed to minimize this impact.
Table 2: Data distribution used for cancer diagnosis. (xdenotes the proportion of positive slides)
@@
Table 3: Data distribution used for cancer diagnosis. (The xrepresents the proportion of each class in the total data volume)
@@
By strategically managing added noise and aggregating updates from multiple centers, FACL effectively maintains commendable model performance while ensuring privacy. In our investigation, we aimed to strike a delicate balance between privacy preservation and model efficacy. We conducted experiments introducing noise at varying magnitudes, discovering that a noise factor of $z=0.1$ provides optimal protection without compromising the overall performance of the model. The culmination of our efforts is encapsulated in the proposed FACL algorithm, integrating contrastive learning and differential privacy, detailed in Algorithm 1. In each iteration, the server disseminates the global model to all participating parties for localized training. Subsequently, the updated local models from each party contribute to refining the global model through a weighted average. We denote the number of rounds of FL training as $K$. Please assume that the WSIs data and its weakly supervised labels are distributed over $M$ independent centers: $(\bm{X},\bm{Y})=\left\{\left\{(\bm{X}_{1,j},Y_{1,j})\right\},...,\left\{(\bm{X}_{M,j},Y_{M,j})\right\}\right\}$, where $\left\{(\bm{X}_{i,j},Y_{i,j})\right\}=\left\{(\bm{X}_{i,1},Y_{i,1}),...,(\bm{X}_{i,N_{i}},Y_{i,N_{i}})\right\}$ denotes the $N_{i}$ pair of WSIs data and the corresponding labels for the $i$th client center, $\bm{X}_{i,j}$ is a single WSI. After data preprocessing and feature extraction, the data are represented as $(\bm{X}^{\prime},\bm{Y})=\left\{\left\{(\bm{X}^{\prime}_{1,j},Y_{1,j})\right\},...,\left\{(\bm{X}^{\prime}_{M,j},Y_{M,j})\right\}\right\}$. The server-side model is $f_{s}$ and the local client’s neural network model $\bm{f}_{center}=\left\{\bm{f}_{1},...,\bm{f}_{M}\right\}$. Each model $\bm{f}_{i}$ contains a projection module $f_{i,pro}$, an attention module $f_{i,att}$, and a prediction layer $f_{i,pre}$. W-Ave: Use the attention scores predicted by all patches as weights to average the feature embeddings of all patches. The parameters of the local model are denoted as $\bm{W}_{1},...,\bm{W}_{M}$, and the parameters of the global model are $\bm{W}_{g}$. ${\rm Opt}_{i}$ is an optimizer that updates the model parameters.
4 Experiments and results
This section introduces the datasets used for prostate cancer diagnosis and Gleason grading tasks. We provide details on data distribution and the experimental setup. Additionally, we assess the effectiveness of FL in both diagnostic and Gleason grading tasks. We compare performance differences between FedAvg and single-center models, as well as between FACL and FedAvg. To enhance interpretability, we employ a heatmap to visualize cancerous regions identified by the model.
@@ 4.1 Datasets
In this study, we delved into the diagnosis of prostate cancer, adopting a two-level classification approach to differentiate between benign and malignant conditions. Furthermore, we examined the Gleason grading of prostate cancer, employing a six-level classification system based on the International Society of Urological Pathology (ISUP bulten2022artificial ) categories: 0, 1, 2, 3, 4, and 5. A series of preprocessing steps were implemented on the datasets to demonstrate the efficacy of the FL model. These steps aimed to enhance the heterogeneity of the datasets by splitting those with a substantial number of samples into multiple independent centers, thus expanding the number of client centers within the FL model. The distribution of datasets utilized in the cancer diagnosis task is presented in Table 2, the DiagSet-B and PANDA datasets have been split into two separate centers, namely DiagSet-B-1, DiagSet-B-2, PANDA-1, and PANDA-2, with the proportion of positive data ranging from 21% to 90%. Similarly, the dataset distribution for Gleason grading is illustrated in Table 3, the PANDA datasets have been split into four separate centers, namely PANDA-1-1, PANDA-1-2, PANDA-2-1, and PANDA-2-2, where the categories of ISUP 0-5 exhibit a notable imbalance. According to the definition of federated learning classification, our work used horizontal federated learning yang2019federated .
For the prostate cancer diagnosis task, we curated datasets from three hospitals (Hebei-1, Hebei-2, Nanchang) and two public sources (DiagSet-B, PANDA) for training purposes. To evaluate the effectiveness of our approach, we used a private hospital dataset (QHD) and a public dataset (DiagSet-A) for testing. Hebei-1 and Hebei-2 are two hospital datasets from Hebei Province, China. Nanchang is a hospital dataset from Nanchang, China. The QHD dataset, sourced from a hospital in Qinhuangdao, China, comprises a total of 765 pathological images, with 433 of them being positive. DiagSet-A, a subset of the DiagSet data, comprises 430 pathological images, with 227 of them being positive.
To assess the accuracy of prostate cancer Gleason grading, we employed two private datasets (Hebei-1, Hebei-2) and one public dataset(PANDA) for training purposes. For the evaluation phase, we employed a private hospital dataset (Nanchang). This allowed us to gauge the performance and reliability of our approach.

### v5

@@ 3 Methods
In this section, we provide an overview of the proposed FACL framework illustrated in Fig. 1. FL retains data locally at individual centers, allowing patch-level segmentation and feature extraction of tissue regions directly from pathological images at the client level. At each client site, the WSIs were cropped into patch images with a size of $224\times 224$ pixels. Subsequently, a Swin transformer, augmented with convolutional operations, was employed to extract discriminative features. These features were then locally preserved as inputs for subsequent model training within the federated learning framework. Each client trains the model by transmitting model parameters to the central server. Subsequently, the server performs weighted averaging by collecting all client model parameters and then distributes the updated parameters to all centers for the next round of training. In addition to these procedures, we enhance dataset protection by incorporating differential privacy. This involves the introduction of random noise during the aggregation of the model parameters at the server.
Initially, each client employs a CTransPath [72] feature extractor to acquire and store features from their local datasets. Subsequently, within the FACL framework, AttMIL [73], a weakly supervised model with an attention mechanism, is incorporated to diagnose and grade gigapixel WSIs in prostate cancer while ensuring patient privacy and enhancing model performance trained using individual client data. To address the non-IID challenge, attention-consistency learning was introduced. Despite the imbalanced data distribution among local centers, FACL can rectify these deviations, enabling the model to achieve optimal results. In general, different datasets exhibit heterogeneity, and federated learning can mitigate heterogeneity challenges. However, severe heterogeneity was observed in the pathological images. Hence, we propose an attention-consistency mechanism. This mechanism enables each client to obtain attention patterns from other clients after the server aggregates the model updates, thereby allowing each client’s model to incorporate diverse attention patterns. This approach enhances the learning capability of individual clients, ultimately leading to more robust performance in the server model.
@@ 3.1 Data preprocessing
High-resolution histopathological imaging often requires time-consuming analyses and is susceptible to interference from complex backgrounds [74]. An integral step in data processing involves cropping the WSIs into smaller sizes suitable for model input. The removal of the background region is crucial because it lacks informative content. Patches ($224\times 224$ pixels) were cropped without overlap in regions with significant tissue presence using the Otsu thresholding method [75]. Subsequently, the features of each patch were extracted using the CTransPath [72] feature extractor and aggregated for consolidated storage across the entire image. Notably, unlike the ImageNet-pre-trained model, the CTransPath extractor was pre-trained on multiple histopathology datasets.
@@ 3.2 Weakly-supervised learning on WSIs
At each local center, we employed a weakly supervised model AttMIL, as detailed in [73]. AttMIL integrates an attention mechanism that forms the basis for FL across local centers. Fig. 2 depicts the attention mechanism pipeline comprising three modules: $f_{pro}$ for projection, $f_{att}$ for attention, and $f_{pre}$ for prediction. We utilize the gating mechanism in conjunction with $\tanh(\cdot)$ nonlinearity. This mechanism enabled the model to dynamically modulate the importance of different parts of the input sequence when making predictions. In this context, we utilized the feature vector extracted from CTransPath as the input for the model. The attention mechanism facilitates the computation of attention scores for each instance, aiding in the identification of similarities or dissimilarities between instances. This approach, particularly relevant in computational pathology, provides interpretability and reasoning for regions of interest, which would be particularly beneficial for medical professionals in the diagnosis of WSIs.
Fig. 2: Specific implementation details of attention mechanism in FACL
We provide detailed layer-by-layer explanations, beginning with the projection module. This module, comprising contiguous, trainable, and fully connected layers, transforms fixed feature embeddings from a pre-trained encoder into a more compact feature space tailored to histopathological images related to the chosen disease model. The initial fully connected layer $\bm{W_{1}}\in\mathbb{R}^{512\times 768}$ compresses 768-dimensional patch-level features $\bm{z}_{k}$ into a 512-dimensional vector $\bm{h}_{k}=\bm{W}_{1}\bm{z}{k}^{\top}$. The resulting set of N patch levels $\bm{h}_{k}$ in the entire slide image is represented by $\bm{H}\in\mathbb{R}^{N\times 512}$.
If we consider the first two attention matrices, $\bm{U}a$ and $\bm{V}a$, both of dimensions $\mathbb{R}^{256\times 512}$, as the shared attention backbone for all classes, the attention network can be divided into n parallel attention branches, denoted as $\bm{W}_{a,1},...,\bm{W}_{a,n}$, each with dimensions $\mathbb{R}^{1\times 256}$. Simultaneously, we created $n$ parallel independent classifiers $\bm{W}_{c,1},...,\bm{W}_{c,n}$ to assess each class-specific slide-level representation. The attention score for the kth patch of the mth class, denoted by $a_{k,m}$ and computed using Equation (1), contributes to the aggregation of the slide-level feature representation. The resulting slide-level feature representation for the mth class is denoted by $h_{slide,m}\in\mathbb{R}^{1\times 512}$ and is computed using Equation (2).
@@
The slide-level score $s_{slide,m}$ is derived from the classification layer $\bm{W}_{c,m}\in\mathbb{R}^{1\times 512}$ using the formula $s_{slide,m}=\bm{W}_{c,m}h_{slide,m}^{\top}$. During the inference, the predicted probability distribution for each class was obtained by applying the SoftMax function to the slide-level prediction scores.
Fig. 3: Attention consistency learning between server and client models. The process unfolds in two main steps during each iteration of FACL training. First, each client is trained on its private data, generating the weight parameters $W_{1}$, $W_{2}$, and $W_{3}$. Second, these weights are transferred to the server model to obtain average weights $W$. The weight parameter $W$ is then replicated as a copy for each client. When each client and its server counterpart were provided the same input for the next training iteration, we imposed an attention-consistency constraint between the generated attention scores. This approach encouraged each client model to emulate the learning abilities of the server model.
@@ 3.3 Federated learning on WSIs
In this section, we elaborate the details of the proposed federated attention-consistent learning method and the differential privacy techniques used to provide additional protection to the model.
3.3.1 Federated attention consistent learning
Inspired by contrastive learning, we propose a federated attention-consistency learning method, as shown in Fig. 3. The FACL comprises two repetitive steps: individual client model training and server model fusion. In the first step, each client was trained using private data. In the second step, the weights from the client models were transferred to update the server model. To ensure that each client model possessed capabilities equivalent to those of the server model, we initially duplicated the server model for each client in the first step. Moreover, when the input images were identical, attention consistency constraints were applied between each client ($W_{1}$, $W_{2}$, and $W_{3}$) and its corresponding server counterpart ($W$), thereby enhancing the performance of each client model. The alignment process for attentional consistency is detailed as follows. In federated learning, the server model collects and synthesizes various attention mechanisms from all the client models. This collective insight is redistributed to each client, allowing them to indirectly benefit from others’ knowledge. Using a copy of the server model, the attention distribution is calculated, which is then considered the baseline or “ground truth” for the subsequent training phase of the clients. Then, the KL loss between the “ground truth” attention distribution and the attention distribution computed by the client model is calculated.
The loss of the local model comprises two components. The first part is a standard loss, such as the cross-entropy loss in supervised learning, denoted by $\mathcal{L}_{CE}$ in this study. The second part is the consistent loss, denoted by $\mathcal{L}_{FACL}$, which measures the disparity between the attention scores of the client- and server-side models. Given the predicted probability value of the classification model as “logits,” the label as “y,” and the attention scores of the local model and the server model as $A_{local}$ and $A_{s}$, respectively, the total loss can be expressed as follows:
@@ 10                   $s_{i,j}=f_{pre}(\bm{h}_{bag_{i,j}})$ ;
                   // Calculate attention consistency
@@ 18             $\bm{W}_{i}\leftarrow\bm{W}_{s}$ ;
Here, $\bm{\mu}$ represents the hyperparameter controlling the weight of attention-consistent loss. $\mathcal{L}_{FACL}$ denotes the Kullback-Leibler divergence [76], which quantifies the similarity between two probability distributions.
@@ 3.3.2 FACL with differential privacy
By embracing the principles of data differential privacy, we seamlessly integrate this approach into the FL server, specifically during the phase in which the server receives parameter updates from individual client models. The crux of differential privacy involves introducing noise into model weights before transmitting them to a centralized server. To achieve this, we utilized a noise generator denoted by $N_{g}(\cdot)$, generating Gaussian random noise $\bm{N}\sim(0,\bm{z}^{2}\eta^{2})$, where $\bm{z}$ indicates the noise level and $\eta$ represents the standard deviation of the weight of the neural network. The FedAvg model with added noise and the FACL model are denoted as FedAvg-$N$ and FACL-$N$, respectively. The noise is meticulously adjusted to prevent a significant deterioration of the learned weights or overall model performance. Achieving an optimal balance entails introducing adequate noise to safeguard privacy while allowing the model to extract valuable information [51, 77]. In FL, noisy weight updates from multiple centers are consolidated on the server. Although the introduction of noise into the weights may have a marginal impact on the learned weights, FL was designed to minimize this effect.
By strategically managing added noise and aggregating updates from multiple centers, the FACL effectively maintains commendable model performance while ensuring privacy. In this investigation, we aimed to strike a delicate balance between privacy preservation and model efficacy. We conducted experiments by introducing noise of varying magnitudes and discovered that a noise factor of $z=0.1$ provides optimal protection without compromising the overall performance of the model. The culmination of our efforts is encapsulated in the proposed FACL algorithm, which integrates attention-consistent learning and differential privacy, detailed in Algorithm 1. In each iteration, the server disseminates the global model to all the participating parties for localized training. Subsequently, the updated local models from each party contribute to refining the global model using a weighted average. The number of FL training rounds is denoted by $K$. Please assume that the WSIs data and its weakly supervised labels are distributed over $M$ independent centers: $(\bm{X},\bm{Y})=\left\{\left\{(\bm{X}_{1,j},Y_{1,j})\right\},...,\left\{(\bm{X}_{M,j},Y_{M,j})\right\}\right\}$, where $\left\{(\bm{X}_{i,j},Y_{i,j})\right\}=\left\{(\bm{X}_{i,1},Y_{i,1}),...,(\bm{X}_{i,N_{i}},Y_{i,N_{i}})\right\}$ denotes the $N_{i}$ pair of WSIs data and the corresponding labels for the $i$th client center, $\bm{X}_{i,j}$ is a single WSI. After data preprocessing and feature extraction, the data are represented as $(\bm{X}^{\prime},\bm{Y})=\left\{\left\{(\bm{X}^{\prime}_{1,j},Y_{1,j})\right\},...,\left\{(\bm{X}^{\prime}_{M,j},Y_{M,j})\right\}\right\}$. The server-side model is $f_{s}$ and the local client’s neural network model $\bm{f}_{center}=\left\{\bm{f}_{1},...,\bm{f}_{M}\right\}$. Each model $\bm{f}_{i}$ contains a projection module $f_{i,pro}$, attention module $f_{i,att}$, and prediction layer $f_{i,pre}$. W-Ave: The attention scores predicted by all patches are used as weights to average the feature embeddings of all patches. The parameters of the local model are denoted by $\bm{W}_{1},...,\bm{W}_{M}$ and those of the global model are $\bm{W}_{g}$. ${\rm Opt}_{i}$ is an optimizer that updates the model parameters.
4 Experiments and results
This section introduces the datasets used for prostate cancer diagnosis and Gleason grading, providing details on the data distribution and experimental setup. We assess the effectiveness of FL in both diagnostic and Gleason grading tasks and compare the performance differences between the FedAvg and single-center models. Additionally, we utilize FedAvg and FedAvg-$N$ as baseline comparisons for FACL and compare them with the results of the FACL and FACL-$N$ models, respectively. To enhance interpretability, a heat map is employed to visualize the cancerous regions identified by the model.
Table 2: Data distribution used for cancer diagnosis (xdenotes the proportion of positive slides)
@@ 4 Experiments and results
Table 3: Data distribution used for cancer diagnosis (The xrepresents the proportion of each class in the total data volume)
@@ 4.1 Datasets
In this study, we investigated the diagnosis of prostate cancer using a two-level classification approach to differentiate between benign and malignant conditions. Furthermore, we examined the Gleason grading of prostate cancer by employing a six-level classification system based on the International Society of Urological Pathology (ISUP [15]) categories 0, 1, 2, 3, 4, and 5. A series of preprocessing steps was implemented on the datasets to demonstrate the efficacy of the FL model. These steps aim to enhance the heterogeneity of the datasets by splitting those with a substantial number of samples into multiple independent centers, thus expanding the number of client centers within the FL model. The distribution of the datasets utilized in the cancer diagnosis task is presented in Table 2. The DiagSet-B and PANDA datasets were split into two separate centers, DiagSet-B-1, DiagSet-B-2, PANDA-1, and PANDA-2, with the proportion of positive data ranging from 21% to 90%. Similarly, the dataset distribution for Gleason grading is illustrated in Table 3, the PANDA datasets have been split into four separate centers, namely PANDA-1-1, PANDA-1-2, PANDA-2-1, and PANDA-2-2, where the categories of ISUP 0-5 exhibit a notable imbalance. According to the definition of federated learning classification, our study uses horizontal federated learning [78].
For the prostate cancer diagnosis task, we curated datasets from three hospitals (Hebei-1, Hebei-2, and Nanchang) and two public sources (DiagSet-B and PANDA) for training. To evaluate the effectiveness of our approach, we tested a private hospital dataset (QHD) and a public dataset (DiagSet-A). Hebei-1 and Hebei-2 represent hospital datasets from the Hebei Province, China, while Nanchang denotes a hospital dataset from Nanchang, China. The QHD dataset, sourced from a hospital in Qinhuangdao, China, comprised 765 pathological images, 433 of which were positive. DiagSet-A, a subset of the DiagSet data, comprised 430 pathological images, of which 227 were positive.
To assess the accuracy of prostate cancer Gleason grading, we employed two private datasets (Hebei-1 and Hebei-2) and one public dataset (PANDA) for training purposes. In the evaluation phase, a private hospital dataset (Nanchang) was utilized. This allowed us to evaluate the performance and reliability of the proposed approach.